OpenStudy (sierra_tibbetts):

Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car. Part A: Write functions to represent Cory and Roger's collections throughout the years. Part B: How many cars does Cory have after 6 years? How many does Roger have after the same number of years? Part C: After approximately how many years is the number of cars that Cory and Roger have the same? NEED HELP ASAP

OpenStudy (dumbcow):

Cory will have an exponential function \[15(1+r)^t\] r = 20% = 0.2 \[15(1.2)^t\] Roger will have a linear function because the increase is constant \[40 + t\] Part C: \[15 (1.2)^t = 40 +t\] Plug in various values for t until both sides are almost equal Notice initially Roger will have more, as t increases Cory will get closer to Roger. If Cory is larger then reduce value of t.

OpenStudy (mrnood):

When you post a question here those who try to help must assume that you have a basic understanding of the topic you are studying. @dumbcow 's answer above is helpful and complete If there are parts of the process you don't understand then maybe you could explain your difficulty and ask for help on a particular point. If you 'don't understand' any of the concepts or points made in that answer then you are probably not ready to tackle this question, and need to go back to your study notes or text and see if you can understand the principles - in this case they are exponential growth and linear equations. Please help those who are trying to help you.